# An Invitation to Deep Reinforcement Learning

### v1

@@ 4.3 Replay Buffers
However, datasets in RL are typically quite biased (Nikishin et al., 2022). Initially, the bias is towards states around the starting region and later towards high reward trajectories, in particular if $\epsilon$ is decayed. Uniformly sampling data from the replay buffer is suboptimal because some samples are more informative (less redundant) than others. Prioritized experience replay (Schaul et al., 2015) addresses this issue by sampling data points proportionally to their loss when they were used the last time. This measures which samples are not well-fitted yet. New samples are given the maximum priority to make sure they are used at least once. Today, prioritized replay buffers are part of the standard RL toolset and used in various Q-learning based methods (Hessel et al., 2018; Badia et al., 2020).
@@ 5.2.1 Long Time Horizons
Optimizing for long time horizons is in general challenging, as the future is typically hard to predict. A common technique to address this is by limiting the effective time horizon of the optimization by applying a discount factor $\gamma\in[0,1]$ to future rewards:
@@
Due to the recursive nature of the Bellman equation, future rewards will at some point implicitly be multiplied by almost 0 when using a discount factor $\gamma<1$. This softly limits the effective time horizon. Doing so biases the objective, but is often used in practice to improve convergence. For clarity, we will omit discount factors from equations, except when presenting concrete algorithms.
@@ 5.2.4 Self-Overfitting
Another way to address the problem of sparse rewards is to maximize a different reward function that provides dense feedback at every time step. Designing this reward function is called reward shaping (Ng et al., 1999). Optimizing for shaped rewards can lead to suboptimal policies wrt. to the original reward function, but in practice often results in better policies because shaped reward functions are designed to ease optimization. When designing a shaped reward function, particular care has to be taken to not introduce “loopholes”, where simple undesired behaviors lead to high rewards (Clark & Amodei, 2016). Neural networks can learn to exploit such loopholes, which is called shortcut learning (Geirhos et al., 2020).
@@ 5.2.5 Deep Networks
In the next section, we will cover Soft Actor-Critic, a popular algorithm based on Q-learning that incorporates many of the ideas discussed in this section into a single algorithm.
@@ 5.3 Soft Actor-Critic (SAC)
The temperature parameter $\alpha$ is a hyperparameter defining the tradeoff between exploration and performance. Besides balancing exploration, a second advantage is that the network is encouraged to put equal probability on equally good actions. As is typical in value learning, the goal is to predict this objective using a Q-function. Soft Actor-Critic trains two $Q$-functions to address the maximization bias, as introduced in Section 5.2.3. It also uses target networks $Q^{\prime}$ as objective to stabilize training, as discussed in Section 5.2.2. The Q-functions are trained with temporal difference learning with the following objective, where $a^{\prime}$ is a sample from the probabilistic policy:
@@ 6 On-policy Reinforcement Learning
We are now going to extend the idea of stochastic policy gradients presented in Section 3.2 to sequential decision making problems. The objective in this setting, the sum of rewards, depends on which actions the policy will predict in future states, which makes computing the policy gradient difficult. Fortunately, it is still possible to compute the policy gradient in sequential settings when the data is collected on-policy. On-policy RL describes the setting where the data collecting policy $\pi_{\beta}$ is the same as the training policy $\pi$. This is achieved by discarding all collected data after every training iteration and collecting new data with the updated policy. On-policy algorithms therefore do not use replay buffers, but store data in temporary buffers that are cleared at the start of the next iteration of the algorithm. In principle, it is also possible to use Q-learning in on-policy settings, but in practice this is rarely done because policy gradient methods empirically achieve better performance in on-policy settings (Mnih et al., 2016). On-policy learning is challenging because there are strong correlations between sequentially collected samples, and rare examples can only be used once. However, on-policy learning with stochastic policy gradients, unlike Q-learning, guarantees convergence to a local minimum also for deep neural networks (Wang et al., 2020b). In the following, we will first introduce the vanilla on-policy policy gradient algorithm REINFORCE. REINFORCE has a number of drawbacks, such as sample inefficiency and high variance. We will hence also discuss ways to mitigate these problems. While we discuss these ideas in the context of policy gradients, some of them can also be applied to value learning methods. Finally, we will introduce proximal policy optimization (PPO), a state-of-the-art algorithm that combines these ideas into a single method.
@@ 6.1 REINFORCE
Here, the factor of proportionality is the average length of an episode (Sutton & Barto, 2018). The policy gradient theorem is a remarkable result because it shows that $\nabla_{\pi}J$ can be computed without backpropagating through the state distribution $d^{\pi}$ or the action-value function $Q^{\pi}$ which makes it feasible to compute this gradient in practice. Additionally, the theorem also applies when $\pi$ is a deep neural network (Wang et al., 2020b). It is important to note that, unlike in value learning, the Q-function in Eq. (28) is the true Q-function, describing the expected return when taking action $a$ in state $s$ and following the policy $\pi$ afterwards. We use parentheses to emphasize that the Q-function is a constant in the gradient computation. When training with on-policy data, the frequency of the actions we observe are also dependent on the probability of the policy taking that action. This can be corrected for by using the log derivative trick $\nabla\log f(x)=\frac{\nabla f(x)}{x}$:
@@
Vanilla REINFORCE policy gradients have high variance because they use Monte Carlo samples to estimate the Q-function and expectation. Revisiting the chess example, if the policy plays a perfect opening but makes a mistake in the middle game and loses, then all opening moves will be down-weighted as well. Discovering which actions contributed to the return is called the credit assignment problem. The objective $G$, here the return $\sum_{i=t}^{N}r_{i}$, is the feedback signal from the environment used to "reinforce" the actions. One way to reduce the variance of the gradient is to normalize the objective $G$ by subtracting a baseline $b(s)$, which can be any constant or function that depends on the state (Williams, 1992; Weaver & Tao, 2001):
@@ 6.2.2 Reward Locality
Discount factors also limit the maximum time horizon that is optimized for, since the contribution of the reward will approach zero for larger $i$. Using discount factors distorts the objective and hence biases the solution. The advantage is that the model may learn faster since actions are only reinforced by the next few rewards. The best value of the discount factor depends on the environment and reward function. While tuning is required in practice, an initial value of 0.99 is often recommended (Andrychowicz et al., 2020). The discount factor can be applied whenever we operate with time sequences. For clarity, we will set the discount factor to 1 in the following.
The assumption of reward locality does not hold in all environments. Consider for example again the game of chess where the reward is +1 for a win and -1 for a loss at the end of the game and 0 otherwise. Using a discount factor in such an environment does not ease learning, as all rewards except for the last one are zero. The only effect of a discount factor here would be that the objective prefers fast wins over wins in games with more steps. Like in off-policy learning, one way to address the problem of sparse rewards is to change the reward function such that it becomes more local and hence denser. This is called reward shaping (Ng et al., 1999). In chess, capturing of a piece could be positively rewarded while losing a piece could be negatively rewarded. The action of capturing a piece now is local with respect to the reward for capturing a piece. However, in chess not all captures are beneficial in the long term. In general, reward shaping biases the solution wrt. the original reward function, but can lead to better policies given a fixed amount of computation because learning can be faster.
@@ 6.2.3 Uncertain Futures
$\lambda=0$ recovers Eq. (37) (we define $10^{0}=$ 1), while $\lambda=1$ recovers the standard Monte Carlo return. This $\lambda$-return objective is also called an eligibility trace. Using a value function to predict expected returns as targets or using discount factors are ways to deal with the uncertainty of returns. Some works try to learn the distribution of returns instead, although this idea is mostly used in the context of Q-learning (Bellemare et al., 2017; Dabney et al., 2018b; a).
@@ 6.3 Proximal Policy Optimization (PPO)
Proximal policy optimization (PPO) (Schulman et al., 2017) is a widely used RL algorithm that combines many extensions to the classic REINFORCE algorithm that we discussed in Section 6.1. PPO has been used successfully to master complex games (Berner et al., 2019), learn autonomous driving planners (Zhang et al., 2021), control drones (Kaufmann et al., 2023) or tune large language models (Ouyang et al., 2022).
@@ Appendix E Related Work
The most widely cited introduction to reinforcement learning is Sutton & Barto (2018) which is a 526-page-long textbook. It puts a strong focus on theoretical foundations and methods using tabular representations or linear function approximation. For such problems, much stronger theoretical guarantees can be obtained than for the non-linear function approximation problems that we considered in this work. The textbook also discusses applications of RL in psychology and neuroscience. As such, Sutton & Barto (2018) is complementary to this work, and we recommend it for readers that want to deeply familiarize themselves with the field. A lot of introductions (Mousavi et al., 2016; Li, 2017; 2018), books (Szepesvári, 2010; François-Lavet et al., 2018; Sutton & Barto, 2018), surveys (Kaelbling et al., 1996; Arulkumaran et al., 2017; Wang et al., 2020a; Lazaridis et al., 2020; Joshi et al., 2021; Wang et al., 2022) and tutorials (Harmon & Harmon, 1996; Li & Pyeatt, 2004; Gosavi, 2009; Levine et al., 2020; Vidyasagar, 2023) have been written over the years on the topic of RL. This work adds to the literature by introducing RL from the alternative angle of optimizing non-differentiable objectives. We introduce the reader to the most important ideas in the field and show the close connection between supervised learning and RL.
/a>

### v2

@@ 4.3 Replay Buffers
However, datasets in RL are typically quite biased (Nikishin et al., 2022). Initially, the bias is towards states around the starting region and later towards high reward trajectories, in particular if $\epsilon$ is decayed. Uniformly sampling data from the replay buffer is suboptimal because some samples are more informative (less redundant) than others. Prioritized experience replay (Schaul et al., 2015) addresses this issue by sampling data points proportionally to their loss when they were used the last time. This measures which samples are not well-fitted yet. New samples are given the maximum priority to make sure they are used at least once. Since prioritized experience replay automatically adjusts the data distribution based on the prediction capabilities of the model it can also be viewed as an automatic curriculum learning method (Portelas et al., 2020). Today, prioritized replay buffers are part of the standard RL toolset and used in various Q-learning based methods (Hessel et al., 2018; Badia et al., 2020).
@@ 5.2.1 Long Time Horizons
Optimizing for long time horizons is in general challenging, as the future is typically hard to predict. The same state and action pair can lead to different returns as the future is typically stochastic. Observed returns can therefore have high variance. A common technique to reduce this variance is by limiting the effective time horizon of the optimization by applying a discount factor $\gamma\in[0,1]$ to future rewards:
@@
Due to the recursive nature of the Bellman equation, future rewards will at some point implicitly be multiplied by almost 0 when using a discount factor $\gamma<1$. This softly limits the effective time horizon. Doing so biases the objective, but is often used in practice to improve convergence. Limiting the effective time horizon with discount factors is a general idea that can be used with almost any RL algorithm. For clarity, we will omit discount factors from equations, except when presenting concrete algorithms.
@@ 5.2.4 Self-Overfitting
N-step returns are a general idea that can be employed with any algorithm that learns a value/Q-function. The hyperparameter $n$ represents a trade-off, where the bias coming from incorrectly estimated target Q-function decreases, and the variance coming from the stochastic nature of future rewards increases, with higher values of $n$. Another way to address the problem of sparse rewards is to maximize a different reward function that provides dense feedback at every time step. Designing this reward function is called reward shaping (Ng et al., 1999). Optimizing for shaped rewards can lead to suboptimal policies wrt. to the original reward function, but in practice often results in better policies because shaped reward functions are designed to ease optimization. Shaping rewards is independent of the learning algorithm and hence can be used with any RL algorithm. When designing a shaped reward function, particular care has to be taken to not introduce “loopholes”, where simple undesired behaviors lead to high rewards (Clark & Amodei, 2016). Neural networks can learn to exploit such loopholes, which is called shortcut learning (Geirhos et al., 2020).
@@ 5.2.5 Deep Networks
Many algorithms exist that extend the basic Deep Q-learning algorithm (Schaul et al., 2015; Hasselt et al., 2016; Fujimoto et al., 2018; Hessel et al., 2018; Kapturowski et al., 2019; Badia et al., 2020; Kapturowski et al., 2023). In the next section, we will cover Soft Actor-Critic as a concrete example. Soft Actor-Critic is a popular algorithm based on Q-learning that incorporates many of the ideas discussed in this section into a single algorithm.
@@ 5.3 Soft Actor-Critic (SAC)
The temperature parameter $\alpha$ is a hyperparameter defining the tradeoff between exploration and performance and $\mathcal{H}(\pi(\cdot|s))=\mathbb{E}_{a\sim\pi}[-\log\pi(a|s)]$. Besides balancing exploration, a second advantage is that the network is encouraged to put equal probability on equally good actions. As is typical in value learning, the goal is to predict this objective using a Q-function. Soft Actor-Critic trains two $Q$-functions to address the maximization bias, as introduced in Section 5.2.3. It also uses target networks $Q^{\prime}$ as objective to stabilize training, as discussed in Section 5.2.2. The Q-functions are trained with temporal difference learning with the following objective, where $a^{\prime}$ is a sample from the probabilistic policy:
@@ 6 On-policy Reinforcement Learning
We are now going to extend the idea of stochastic policy gradients presented in Section 3.2 to sequential decision making problems. The objective in this setting, the sum of rewards, depends on which actions the policy will predict in future states, which makes computing the policy gradient difficult. Fortunately, it is still possible to compute the policy gradient in sequential settings when the data is collected on-policy. On-policy RL describes the setting where the data collecting policy $\pi_{\beta}$ is the same as the training policy $\pi$. This is achieved by discarding all collected data after every training iteration and collecting new data with the updated policy. On-policy algorithms therefore do not use replay buffers, but store data in temporary buffers that are cleared at the start of the next iteration of the algorithm. In principle, it is also possible to use Q-learning in on-policy settings, but in practice this is rarely done because policy gradient methods empirically achieve better performance in on-policy settings (Mnih et al., 2016). On-policy learning is challenging because there are strong correlations between sequentially collected samples, and rare examples can only be used once. However, on-policy learning with stochastic policy gradients, unlike Q-learning, guarantees convergence to a local minimum also for deep neural networks (Wang et al., 2020b). In the following, we will first introduce the vanilla on-policy policy gradient algorithm REINFORCE. REINFORCE has a number of drawbacks, such as sample inefficiency and high variance. We will hence also discuss ways to mitigate these problems. Finally, we will introduce proximal policy optimization (PPO), a state-of-the-art algorithm that combines these ideas into a single method.
@@ 6.1 REINFORCE
Here, the factor of proportionality is the average length of an episode (Sutton & Barto, 2018). For a derivation of the policy gradient theorem, see Sutton & Barto (2018) or Levine (2023). The policy gradient theorem is a remarkable result because it shows that $\nabla_{\pi}J$ can be computed without backpropagating through the state distribution $d^{\pi}$ or the action-value function $Q^{\pi}$ which makes it feasible to compute this gradient in practice. Additionally, the theorem also applies when $\pi$ is a deep neural network (Wang et al., 2020b). It is important to note that, unlike in value learning, the Q-function in Eq. (28) is the true Q-function, describing the expected return when taking action $a$ in state $s$ and following the policy $\pi$ afterwards. We use parentheses to emphasize that the Q-function is a constant in the gradient computation. When training with on-policy data, the frequency of the actions we observe are also dependent on the probability of the policy taking that action. This can be corrected for by using the log derivative trick $\nabla\log f(x)=\frac{\nabla f(x)}{x}$:
@@
Vanilla REINFORCE policy gradients have high variance because they use Monte Carlo samples to estimate the Q-function and expectation. Revisiting the chess example, if the policy plays a perfect opening but makes a mistake in the middle game and loses, then all opening moves will be down-weighted as well. Discovering which actions contributed to the return is called the credit assignment problem. Credit assignment techniques, while not strictly necessary, often significantly simplify optimization. The objective $G$, here the return $\sum_{i=t}^{N}r_{i}$, is the feedback signal from the environment used to "reinforce" the actions. One way to reduce the variance of the gradient is to normalize the objective $G$ by subtracting a baseline $b(s)$, which can be any constant or function that depends on the state (Williams, 1992; Weaver & Tao, 2001):
@@ 6.2.2 Reward Locality
Discount factors also limit the maximum time horizon that is optimized for, since the contribution of the reward will approach zero for larger $i$. Using discount factors distorts the objective and hence biases the solution. The advantage is that the model may learn faster since actions are only reinforced by the next few rewards. The best value of the discount factor depends on the environment and reward function. While tuning is required in practice, an initial value of 0.99 is often recommended (Andrychowicz et al., 2020). Again, the discount factor is a general idea that can be combined with any RL algorithm that addresses multistep problems. For clarity, we will set the discount factor to 1 in the following.
The assumption of reward locality does not hold in all environments. Consider for example again the game of chess where the reward is +1 for a win and -1 for a loss at the end of the game and 0 otherwise. Using a discount factor in such an environment does not ease learning, as all rewards except for the last one are zero. The only effect of a discount factor here would be that the objective prefers fast wins over wins in games with more steps. Like in off-policy learning, one way to address the problem of sparse rewards is to change the reward function such that it becomes more local and hence denser. This is called reward shaping (Ng et al., 1999). In chess, capturing of a piece could be positively rewarded while losing a piece could be negatively rewarded. The action of capturing a piece now is local with respect to the reward for capturing a piece. However, in chess not all captures are beneficial in the long term. In general, reward shaping biases the solution wrt. the original reward function, but can lead to better policies given a fixed amount of computation because learning can be faster. Like discount factors, reward shaping can be used in combination with any RL method.
@@ 6.2.3 Uncertain Futures
$\lambda=0$ recovers Eq. (37) (we define $10^{0}=$ 1), while $\lambda=1$ recovers the standard Monte Carlo return. This $\lambda$-return objective is also called an eligibility trace. Both $n$-step returns and eligibility traces are not specific to policy gradient methods but can be used with any method that learns a value function. Using a value function to predict expected returns as targets or using discount factors are ways to deal with the uncertainty of returns. Some works try to learn the distribution of returns instead, although this idea is mostly used in the context of Q-learning (Bellemare et al., 2017; Dabney et al., 2018b; a).
@@ 6.3 Proximal Policy Optimization (PPO)
Various methods exist that build upon the basic policy gradient idea (Mnih et al., 2016; Espeholt et al., 2018; Petrenko et al., 2020; Cobbe et al., 2021). In this section, we cover a concrete example, proximal policy optimization (PPO) (Schulman et al., 2017), which is a widely used RL algorithm that combines many extensions to the classic REINFORCE algorithm that we discussed in Section 6.1. PPO has been used successfully to master complex games (Berner et al., 2019), learn autonomous driving planners (Zhang et al., 2021), control drones (Kaufmann et al., 2023) or tune large language models (Ouyang et al., 2022).
@@ Appendix E Related Work
Many past RL tutorials (Harmon & Harmon, 1996; Li & Pyeatt, 2004; Gosavi, 2009) do not cover deep RL, since this development happened after these works were published. Mousavi et al. (2016) gives an overview over the deep RL literature, but is shallow in terms of technical details and largely neglects the important topic of policy gradient methods. Levine et al. (2020) focuses on offline RL, whereas we focus on more mature off-policy and on-policy deep RL techniques. As such, Levine et al. (2020) is complementary to our work. Vidyasagar (2023) focuses on RL theory and does not cover actor-critic methods or modern RL algorithms, whereas we cover modern deep RL.
Surveys on RL (Kaelbling et al., 1996; Arulkumaran et al., 2017; Wang et al., 2020a; Lazaridis et al., 2020; Joshi et al., 2021; Wang et al., 2022) typically review the latest research advancements in the field. Our tutorial instead covers ideas that stood the test of time and is as such more suitable as introductory material.
Existing RL Books (Szepesvári, 2010; François-Lavet et al., 2018; Li, 2018; Sutton & Barto, 2018; Lu et al., 2023) (as well as lectures (Silver, 2015; Levine, 2023)) cover a broad range of topics but require a large time commitment to consume. For example, the most widely cited introduction to reinforcement learning is Sutton & Barto (2018) which is a 526-page-long textbook. It puts a strong focus on theoretical foundations and methods using tabular representations or linear function approximation. For such problems, much stronger theoretical guarantees can be obtained than for the non-linear function approximation problems that we considered in this work. The textbook also discusses applications of RL in psychology and neuroscience. As such, Sutton & Barto (2018) is complementary to this work, and we recommend it for readers that have an interest in these topics. François-Lavet et al. (2018) is the closest related manuscript, and can perhaps be seen as a representative of the traditional way to introduce deep reinforcement learning. It introduces deep reinforcement learning as techniques for sequential decision making via the theoretical framework of Markov decision processes. On the contrary, we introduce deep reinforcement learning as a generalization of supervised learning to non-differentiable objectives, which provides an alternative introduction which is more suitable to the larger supervised learning audience. François-Lavet et al. (2018) covers a broad range of ideas in reinforcement learning, giving readers a broad overview about many ideas in the field. However, due to covering so many topics it often lacks the necessary depth for the reader to fully understand the presented ideas, requiring the reader to seek out additional material. Instead, our manuscript focuses on depth, by zooming in on the most important ideas covering them in sufficient detail such that the reader can fully understand the ideas and concrete state-of-the-art implementations of them without conducting additional material. As a result, our manuscript is more than 2x shorter and more suitable for readers that want to apply popular reinforcement learning ideas to domains outside of reinforcement learning, such as robotics, computer vision or generative AI.